OpenStudy (anonymous):

I Medal ! Bert is planning to open a savings account that earns 1.6% simple interest yearly. He wants to earn exactly $144 in interest after 3 years. How much money should he deposit? $2,304 $300 $3,000 $30

OpenStudy (mathmath333):

do u know simple interest formula

OpenStudy (mathmath333):

\(\tt \color{black}{\text{u are given interset ,time and rate u need to find the principal here}}\) \(\large\tt \color{black}{I=Prt}\) \(\large\tt \color{black}{144=P\times 1.6\% \times3}\) \(\large\tt \color{black}{P=\dfrac{144\times 100}{1.6\times 3}}\) \(\large\tt \color{black}{P=3000}\)